Mr. Morton invested $1400, part at 5% and the rest at 8%. His total annual income from these investments was $100. Find the amount he
invested at each rate.

Please help me solve this. I know this much:
X=amount at 5%
Y=amount at 8%
X+Y=100

Answer by hamsanash1981@gmail.com(151)


x = amount at 5%
y = amount at 8%
then X + Y = 1400 ---1
Let the number of years be 1.
simple interest of X = (X*1*5)/100 => 5x/100
Simple interest of Y = (Y*1*8)/100 => 8y/100
Given that his total annual returns from these investment was 100 then,
5x/100 + 8y/100 = 100 ---2
Multiplying equation 2 by 100
5x + 8y = 10000 ---3
Solving equ 1 and 3
8x + 8y = 11200
5x + 8y = 10000
----------------
3x = 1200
x = 400
then Y = 1400 - 400
Y = 1000
therefore 400 invested at 5% and 1000 invested at 8%.